2008 | 2009 | ||
October | January | April | July |
November | February | May | August |
December | March | June | September |
01/14/09
The BM2 magnet was connected to the BM4 PS. Meanwhile, BM2 power has been fixed. According to Leonard, knowing the BLIP beam energy is important. Therefore, a job request to switch back to BM2 has been placed (Paul S.). Checking BM4 PS at the first opportunity is highly desirable (it might have the same problem as BM2).
RHIC access re-certification lockout ("rollout") is scheduled for Friday, Jan. 16th. It will be done in two sessions: 4:00-9:00 am and 4:30 - 9:00 pm.
Linac startup was delayed. There were a few problems with delivering high power beam to BLIP:
An aperture in the low energy beam line did not let matching to the linac. Removed.
Quad triplet between Tank 1 and 2 did not have enough power. DC PS were substituted for the pulsed ones.
One of the RF windows failed. It will be changed during one of maintenance days. With one window, RF power is insufficient to fully bunch the high intensity beam. Currently, the beam current after tank 2 is 29 mA vs 40 required (see Time meeting 01/06/09). This should not affect the polarized beam operations because less power is needed.
Bending magnet 2 power supply failure (no communication to the safety interlock system.) This has been known before Christmas. The problem still existed on Jan. 6, 2009. After the issue of affecting the RHIC startup schedule was brought up, additional people/efforts to fix the magnet were dispatched/planned.
RHIC startup schedule milestones have been shows at the Time meeting today, added to the scheduling physicist webpage. The schedule milestones also have been reported at several management meeting in November.
BtA/AGS studies are wrapping up tonight. Processed data / results are still to see. NSRL running without major interruptions. Because there will be only NSRL running beginning from today, there will be no regular time meetings beginning from today.
BLIP wants to start operations on 5 Jan as scheduled. First beam to Booster on Jan. 10-12 (3.5 weeks before the Blue ring is cold).
Mei has sent preliminary Run 9 pp 500 GeV Start-up plan. We are waiting for detector planning meetings to start discussing the schedule. The first meeting is planned on 11/18/08.
Vladimir L. suggested to move the experiment approval committee meeting from Friday to Wednesday. This will clarify the experiment status earlier and let PS group personnel to prepare for turning on AGS.
Follow-up: AGS studies have been approved. Thus, 2 weeks of BtA + 1 week of AGS studies
K. Brown: BtA studiers meet each day in the MCR conference room at 4pm. We (and NSRL) plan to be running with a 2.5 sec rep rate for protons. Dana has given us a proton limit and Nick K. has put together a beam counter to keep us on track. Activity was limited to just going to DH2&3 (kept off) and cannot exceed 1e14 dumped beam in DH2&3 to keep dirt/water activation under control. Originally, it was thought Siemems would be available and we would primarily dump at either Booster injection or into J10. The evening work consisted of working on having beam on Booster U4 (for BtA studies) ppm with beam on U2 (for NSRL). Check off lists were completed by 7 pm.
10/09/08: The journal has been created.